OpenStudy (anonymous):

At what point on the curve y = 9 + 2ex − 6x is the tangent line parallel to the line 6x − y = 1?

OpenStudy (loser66):

the curve is y = 9 + 2e^x -6x??

OpenStudy (anonymous):

Sorry, yes the curve is y = 9 + 2e^x - 6x

OpenStudy (loser66):

I got the point is (ln3 , 3-6ln3) . Since it's weird, I am not sure.

OpenStudy (anonymous):

Thanks for the help, but I don't think this is the correct point..

OpenStudy (anonymous):

what's the slope of the line? can you calculate it?

OpenStudy (anonymous):

the slope is 6

OpenStudy (anonymous):

what's the slope of the function, i.e., the derivative? find the x that makes them the same.

OpenStudy (anonymous):

I get (x, y) = (ln3, 3)

OpenStudy (anonymous):

f'(x) = 2e^x - 6 = 6 => 2e^x=12 => e^x = 6 => x = ln6

OpenStudy (anonymous):

(ln6, 21-6ln6)

OpenStudy (anonymous):

do you follow?

OpenStudy (anonymous):

yes, that should be the correct one. Thanks!

OpenStudy (anonymous):

do you understand how to do those?

OpenStudy (anonymous):

I have to do more reading to understand it better how to find the y value

OpenStudy (anonymous):

once you have the x that makes the slope you need, you just plug that into the original equation to find y. y=f(ln6) = 9 +2e^(ln6)-6(ln6) => y=9+2(6)-6ln6=9+12-6ln6 = 21-6ln6